OpenStudy (anonymous):

can someone help prove the identitiy for 1+sin(theta)/sin (theta) + cot (theta) -cos(theta)/cos (theta)= 2/sin(theta)

OpenStudy (kc_kennylau):

LaTeX: \[\frac{1+sin\theta}{sin\theta}+\frac{cot\theta-cos\theta}{cos\theta}=\frac2{sin\theta}\]

OpenStudy (anonymous):

need to prove the identity

OpenStudy (kc_kennylau):

just clarifying :)

OpenStudy (anonymous):

thank you

OpenStudy (kc_kennylau):

\[\frac{1+\sin\theta}{\sin\theta}+\frac{\cot\theta-\cos\theta}{\cos\theta}=\frac1{\sin\theta}+1+\frac{\cot\theta}{\cos\theta}-1=\frac1{\sin\theta}+\frac1{\sin\theta}=\frac2{\sin\theta}\]

OpenStudy (anonymous):

Thank you just one question cot(theta)/cos(theta) = 1/sin(theta)

OpenStudy (kc_kennylau):

\[\frac{\cot\theta}{\cos\theta}=\frac{\frac{\cos\theta}{\sin\theta}}{\cos\theta}=\frac1{\sin\theta}\]

OpenStudy (anonymous):

got it. thank you

OpenStudy (kc_kennylau):

no problem :)